OpenStudy (anonymous):

Simplify (4x2+ 12x - 9) + (7x2- x + 3) 11x2+ 12x - 6 11x2+ 13x + 6 11x2+ 11x + 6 11x2+ 11x - 6 Im not sure how to simplify polynomials, if someone could just walk me through it, im sure its not even that hard.. i have a basic idea..

OpenStudy (anonymous):

The parentheses aren't important here because you have nothing to distribute among them. If it were (4x2+ 12x - 9) *minus* (7x2- x + 3) it would be different, but you can just get rid of them here. 4x^2 + 12x - 9 + 7x^2 - x + 3 Then you "combine like terms," that is, you combine terms of the same degree and same variable. 11x^2 +11x - 6 You are not allowed to combine terms of different degrees or different variables because you can't predict what their value will be!

OpenStudy (anonymous):

OHH i see how you get the 11 now

OpenStudy (anonymous):

4x^2 + 7x^2 = 11x^@ 12x - x = 11x -9 + 3 = 11

OpenStudy (anonymous):

Whoops, 11x^2

OpenStudy (anonymous):

Yeah, it's actually a use of the distributive property if you look very carefully. 4x^2 + 7x^2 = x^2(4 + 7) = x^2(11) = 11x^2

OpenStudy (anonymous):

isnt -9 + 3 = -6 ?

OpenStudy (anonymous):

Yeah, sorry, I typed that response very quickly. My original response was correct. -9 + 3 is definitely -6.

OpenStudy (anonymous):

+11x - 6 so the -6 was put there as a minus 6?

OpenStudy (anonymous):

for that part of the equation

OpenStudy (anonymous):

Yeah, so since 4x^2 + 7x^2 = 11x^2 12x - x = 11x -9 + 3 = -6 You can add all those results up to get (11x^2) + (11x) + (-6) and simplify to 11x^2 + 11x - 6.

OpenStudy (anonymous):

thank you, ( I might have a few more, but this one helps alot of them.)

OpenStudy (anonymous):

Yeah, no problem! Math gets a lot easier as you memorize these little things.